# Linear Discriminant Regularized Regression

### v1

@@ A.3 Proof of Theorem 3
By combining (A.18), (A.19), (A.20) and (A.21) and taking the union bounds over $k,\ell\in[L]$, we conclude that
$\|\widehat{H}-H\|_{\infty,1}$ $\lesssim~{}\sqrt{\log(n)\over nL}+\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{L\log(n)\over n}$
$\qquad+\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}\sqrt{\Delta_{\infty}}+L\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}^{2}$
with probability at least $1-n^{-1}$. Invoking Theorem 2 and using (3.3) to simplify the expression complete the proof. ∎
A.4 Proof of Theorem 4
Proof.
Pick any $\ell\in[L]$. Start with
$|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\leq|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}|+|(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+2|\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right)|.$
We work on $\mathcal{E}_{\pi}$ in (A.1) throughout the rest of the proof. For the first term in $\widehat{G}_{\ell}(X)-G_{\ell}(X)$, invoking Lemma 16 with $k=\ell$ and $v=B^{*}_{\ell}$, using the fact that $B_{\ell}^{*\top}\Sigma_{w}B_{\ell}^{*}=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}=\Delta_{\ell\ell}$ and taking a union bounds over $\ell\in[L]$ gives that
$\mathbb{P}\left\{\left\{\max_{\ell\in[L]}|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}|\leq\sqrt{\Delta_{\infty}}\sqrt{aL\log(n)\over n}\right\}\cap\mathcal{E}_{\pi}\right\}\geq 1-2n^{-a}.$ (A.22)
To bound from above
using Cauchy-Schwarz inequality gives
$|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\leq\|\mu_{\ell}^{\top}\Sigma_{w}^{-1/2}\|_{2}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\leq\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}}.$ (A.23)
On the other hand, by using the same arguments of proving (A.3), we have that on the event $\mathcal{E}_{H}$
$|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\leq|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}|+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B(\widehat{\Omega}_{\ell}-\Omega_{\ell})|$
$\leq\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}\|\widehat{B}-B\|_{1,\infty}\|\widehat{\Omega}_{\ell}\|_{1}+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H\widehat{\Omega}-\mathbb{I}_{L})_{\ell}|$
$\leq\left(\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}\|\widehat{B}-B\|_{1,\infty}+\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{\infty}\delta\right)2\|\Omega_{\ell}\|_{1}.$
Here $\delta$ is defined in (A.13). Invoking (A.15), (A.22), (A.23) and Lemma 16 with $v={\bm{e}}_{j}$ and union bounds over $j\in[p]$ yields that, on $\mathcal{E}_{H}$ defined in (A.14),
$|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $~{}\lesssim L\Delta_{\infty}\sqrt{\|\Sigma_{w}\|_{\infty}}\sqrt{L\log(n\vee p)\over n}\|\widehat{B}-B\|_{1,\infty}$
$\qquad+L\delta\sqrt{\Delta_{\infty}^{3}}\sqrt{L\log(n)\over n}$
with probability at least $1-4(n\vee p)^{-2}$. Therefore, with probability at least $1-5(n\vee p)^{-2}$,
$\begin{split}|\widehat{\mu}_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|&~{}\lesssim~{}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}}+\delta\sqrt{\Delta_{\infty}^{3}}\sqrt{L^{3}\log(n)\over n}\\ &\qquad+\Delta_{\infty}\sqrt{L^{3}\log(n\vee p)\over n}\|\widehat{B}-B\|_{1,\infty}.\end{split}$ (A.24)
Finally, conditioning on $Y={\bm{e}}_{k}$, the Gaussian tail of $X\mid Y,{\mathbb{X}},{\mathbb{Y}}$ gives that, for all $t\geq 0$,
$\mathbb{P}\left\{|X^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\geq|\mu_{k}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+t\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}~{}\Big{|}~{}Y={\bm{e}}_{k},{\mathbb{X}},{\mathbb{Y}}\right\}\leq 2e^{-t^{2}/2}.$
Choosing $t=\sqrt{4a\log(n)}$, taking the union bounds over $\ell\in[L]$ and unconditioning yield
$\max_{\ell\in[L]}|X^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\geq\max_{k,\ell\in[L]}|\mu_{k}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+\max_{\ell\in[L]}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{4a\log(n)}$
$\geq 3\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{2,\infty}\sqrt{(a\vee\Delta_{\infty})\log(n)}$ (A.25)
with probability less than $2n^{-a}$. The last step uses (A.23). Collecting (A.2), (A.22), (A.24) and (A.4) and invoking Theorem 3 complete the proof. ∎
A.5 Proof of Theorem 5
For any $x\in\mathbb{R}^{p}$, write
For any $t\geq 0$, define the event
$\mathcal{E}_{t}=\bigcap_{\ell\in[L]}\left\{\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|\leq t~{}\Big{|}~{}{\mathbb{X}},{\mathbb{Y}}\right\}.$ (A.26)
Proof.
By definition, we start with
$\quad\mathcal{R}(\widehat{g})$
$=\sum_{k\in[L]}\pi_{k}\Bigl{\{}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)\neq k\}\mid Y={\bm{e}}_{k}\right]-\mathbb{E}\left[\mathbbm{1}\{g^{*}(X)\neq k\}\mid Y={\bm{e}}_{k}\right]\Bigr{\}}$
$=\sum_{k\in[L]}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)\neq k,g^{*}(X)=k\}\mid Y={\bm{e}}_{k}\right]-\sum_{k\in[L]}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=k,g^{*}(X)\neq k\}\mid Y={\bm{e}}_{k}\right]$
$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y={\bm{e}}_{k}\right]-\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=k,g^{*}(X)=\ell\}\mid Y={\bm{e}}_{k}\right]$
$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\Bigl{\{}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y={\bm{e}}_{k}\right]-\pi_{\ell}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid L=\ell\right]\Bigr{\}}.$
[…]
where, for all $k,\ell\in[L]$,
Observe that the event $\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\cap\mathcal{E}_{t}$ implies
$00>G_{z}^{(\ell|k)}(z)\overset{\mathcal{E}_{t}}{\geq}{1\over 2}\left(\widehat{G}_{k}(X)-\widehat{G}_{\ell}(X)\right)-t\geq-t,\quad\forall~{}t>$ 0 .
By the basic inequality $1+z\leq\exp(z)$ for all $z\in\mathbb{R}$, we obtain that, for any $t>0$,
$\mathcal{R}(\widehat{g})$ $\leq\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\left\{t\pi_{k}\mathbb{E}\left[\mathbbm{1}\{-t\leq G^{(\ell|k)}(X)\leq 0,\widehat{g}(X)=\ell\mid Y={\bm{e}}_{k}\}\right]+\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y={\bm{e}}_{k})\right\}$
$\leq t\sum_{k\in[L]}\pi_{k}\left\{\mathbb{P}\left(-t\leq\max_{\ell\in[L]\setminus\{k\}}G^{(\ell|k)}(X)\leq 0\mid Y={\bm{e}}_{k}\right)+(m-1)\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y={\bm{e}}_{k})\right\}$
$\leq t(m-1)\sum_{k\in[L]}\pi_{k}\max_{\ell\in[L]\setminus\{k\}}\mathbb{P}\left\{-t\leq G^{(\ell|k)}(X)\leq 0\mid Y={\bm{e}}_{k}\right\}+(m-1)\mathbb{P}(\mathcal{E}_{t}^{c})$
Note that
$G^{(\ell|k)}(X)\mid Y={\bm{e}}_{k}~{}\sim~{}N\left(-{1\over 2}\|\mu_{k}-\mu_{\ell}\|_{\Sigma_{w}}^{2}+\log\left(\pi_{\ell}\over\pi_{k}\right),~{}\|\mu_{k}-\mu_{\ell}\|_{\Sigma_{w}}^{2}\right).$
By the mean-value theorem, we conclude
$\mathcal{R}(\widehat{g})$ $\leq(m-1)t^{2}+(m-1)\mathbb{P}(\mathcal{E}_{t}^{c}).$
Choosing
$t=\Delta_{\infty}^{3}\sqrt{sL\log(n\vee p)\log(n)\over n}$
and invoking Theorem 4 complete the proof. ∎
Appendix B Technical lemmas
@@
$\Omega:=D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}.$ (B.1)
@@
${\mathbb{M}}^{\top}B={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}~{}\Omega^{-1},$ (B.2)
$H=D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B=\Omega^{-1},$ (B.3)
$B^{\top}\Sigma_{w}B=\Omega^{-1}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}~{}\Omega^{-1}.$ (B.4)
@@
$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}\leq\pi_{\ell}\left(1\wedge\pi_{\ell}\Delta_{\ell\ell}\right),$ (B.5)
$B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}\leq\left(1\wedge{\pi_{\ell}^{2}\over\pi_{\min}}\Delta_{\ell\ell}\right),$ (B.6)
$\left|\mu_{k}^{\top}B_{\ell}\right|\leq\left(1\wedge\pi_{\ell}\sqrt{\Delta_{kk}\Delta_{\ell\ell}}\right).$ (B.7)
@@
$B=B^{*}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)=\Sigma_{w}^{-1}{\mathbb{M}}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)$
@@
${\mathbb{M}}^{\top}B={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}-{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}B.$
@@
${\mathbb{M}}^{\top}B$ $=\left(\mathbb{I}_{L}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}\right)^{-1}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}$
$=\mathbb{I}_{L}-\left(\mathbb{I}_{L}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}\right)^{-1}$
$=\mathbb{I}_{L}-D_{\pi}^{-1}\left(D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\right)^{-1}$ (B.8)
$={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\left(D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\right)^{-1}$
proving (B.2). By using the above identity (B.8), we have
$D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B=\left(D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\right)^{-1},$
proving (B.3). Since Lemma 1 and (B.3) imply
@@
(B.4) follows from
$B^{\top}\Sigma_{w}B=\Omega^{-1}B^{*\top}\Sigma_{w}B^{*}\Omega^{-1}=\Omega^{-1}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\Omega^{-1}.$
To prove (B.5) – (B.7), pick any $k,\ell\in[L]$. By using (2.6) twice and (2.5), we find that
$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}=\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}\leq\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$ (B.9)
yielding the second bound in (B.5). The other bound in (B.5) follows by observing that
@@
For (B.6), on the one hand, similar arguments yield
$B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}\mu_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}.$
On the other hand, using (B.2) and (B.3) proves
$B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}\leq{\bm{e}}_{\ell}^{\top}\Omega^{-1}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{2}\Omega^{-1}{\bm{e}}_{\ell}\leq 1.$
@@
The following lemmas contain deviation inequalities related with $\widehat{\mathbb{M}}-{\mathbb{M}}$ and $\widehat{\pi}-\pi$.
@@
$=v^{\top}\sum_{k=1}^{L}{1\over n}\sum_{i=1}^{n}\left(X_{i}-{\mathbb{M}}Y_{i}\right)u^{\top}Y_{i}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}$
$={1\over n}\sum_{i=1}^{n}v^{\top}\left(X_{i}-{\mathbb{M}}Y_{i}\right)u^{\top}Y_{i}.$
For $i\in[n]$, by conditioning on ${\mathbb{Y}}$, we know that $W_{i}:=X_{i}-{\mathbb{M}}Y_{i}$ are i.i.d. from $\mathcal{N}_{p}(0,\Sigma_{w})$, hence
${1\over n}\sum_{i=1}^{n}v^{\top}W_{i}u^{\top}Y_{i}~{}\Big{|}~{}{\mathbb{Y}}\sim\mathcal{N}\left(0,{v^{\top}\Sigma_{w}v\over n}{u^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}u\over n}\right).$
@@
$={1\over n}\sum_{i=1}^{n}\left[(X_{i}-{\mathbb{M}}Y_{i})(X_{i}-{\mathbb{M}}Y_{i})^{\top}-\Sigma_{w}\right].$
Recall that, conditioning ${\mathbb{Y}}$, $W_{i}=X_{i}-{\mathbb{M}}Y_{i}$ for $i\in[n]$ are independent, $\mathcal{N}_{p}(0,\Sigma_{w})$. The result follows from standard concentration inequalities for the quadratic term of Gaussian random vectors. ∎
B.1 Concentration inequalities related with $\widehat{\Sigma}-\Sigma$
Recall that $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$ and the definition of the event $\mathcal{E}_{\pi}$ is given in (A.1).

### v2

@@ B.7 Proof of Theorem 8: the reduced-rank estimator of $B$
Recall that $|\mathcal{N}_{p}(1/3)|\leq 7^{p}$ and $|\mathcal{N}_{L}(1/3)|\leq 7^{L}$. Choosing $t=C(p+L)\log(n)$, taking the union bounds over $\mathcal{N}_{p}(1/3)$ and $\mathcal{N}_{L}(1/3)$ and combining with the bound of I complete the proof. ∎
Appendix C Technical lemmas
@@
$\Omega:=D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}.$ (C.1)
@@
${M}^{\top}B={M}^{\top}\Sigma_{w}^{-1}{M}~{}\Omega^{-1},$ (C.2)
$H=D_{\pi}-D_{\pi}{M}^{\top}B=\Omega^{-1},$ (C.3)
$B^{\top}\Sigma_{w}B=\Omega^{-1}{M}^{\top}\Sigma_{w}^{-1}{M}~{}\Omega^{-1}.$ (C.4)
@@
$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}\leq\pi_{\ell}\left(1\wedge\pi_{\ell}\Delta_{\ell\ell}\right),$ (C.5)
$B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq\left(1\wedge{\pi_{\ell}^{2}\over\pi_{\min}}\Delta_{\ell\ell}\right),$ (C.6)
$\left|\mu_{k}^{\top}B_{\ell}\right|\leq\left(1\wedge\pi_{\ell}\sqrt{\Delta_{kk}\Delta_{\ell\ell}}\right).$ (C.7)
@@
$B=B^{*}\left(D_{\pi}-D_{\pi}{M}^{\top}B\right)=\Sigma_{w}^{-1}{M}\left(D_{\pi}-D_{\pi}{M}^{\top}B\right)$
@@
${M}^{\top}B={M}^{\top}\Sigma_{w}^{-1}{M}\left(D_{\pi}-D_{\pi}{M}^{\top}B\right)={M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}-{M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}{M}^{\top}B.$
@@
${M}^{\top}B$ $=\left(\mathbf{I}_{L}+{M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}\right)^{-1}{M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}$
$=\mathbf{I}_{L}-\left(\mathbf{I}_{L}+{M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}\right)^{-1}$
$=\mathbf{I}_{L}-D_{\pi}^{-1}\left(D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}\right)^{-1}$ (C.8)
$={M}^{\top}\Sigma_{w}^{-1}{M}\left(D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}\right)^{-1}$
proving (C.2). By using the above identity (C.8), we have
$D_{\pi}-D_{\pi}{M}^{\top}B=\left(D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}\right)^{-1},$
proving (C.3). Since Lemma 1 and (C.3) imply
@@
(C.4) follows from
$B^{\top}\Sigma_{w}B=\Omega^{-1}B^{*\top}\Sigma_{w}B^{*}\Omega^{-1}=\Omega^{-1}{M}^{\top}\Sigma_{w}^{-1}{M}\Omega^{-1}.$
To prove (C.5) – (C.7), pick any $k,\ell\in[L]$. By using (2.6) twice and (2.5), we find that
$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}=\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}\leq\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$ (C.9)
yielding the second bound in (C.5). The other bound in (C.5) follows by observing that
@@
For (C.6), on the one hand, similar arguments yield
$B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}{M}D_{\pi}{M}^{\top}\Sigma^{-1}\mu_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}.$
On the other hand, using (C.2) and (C.3) proves
$B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq{\bm{e}}_{\ell}^{\top}\Omega^{-1}({M}^{\top}\Sigma_{w}^{-1}{M})^{2}\Omega^{-1}{\bm{e}}_{\ell}\leq 1.$
@@
The following lemmas contain deviation inequalities related with $\widehat{M}-{M}$ and $\widehat{\pi}-\pi$.
@@
$=v^{\top}\sum_{k=1}^{L}{1\over n}\sum_{i=1}^{n}\left(X_{i}-{M}Y_{i}\right)u^{\top}Y_{i}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}$
$={1\over n}\sum_{i=1}^{n}v^{\top}\left(X_{i}-{M}Y_{i}\right)u^{\top}Y_{i}.$
For $i\in[n]$, by conditioning on ${\mathbf{Y}}$, we know that $W_{i}:=X_{i}-{M}Y_{i}$ are i.i.d. from $\mathcal{N}_{p}(0,\Sigma_{w})$, hence
${1\over n}\sum_{i=1}^{n}v^{\top}W_{i}u^{\top}Y_{i}~{}\Big{|}~{}{\mathbf{Y}}\sim\mathcal{N}\left(0,{v^{\top}\Sigma_{w}v\over n}{u^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}u\over n}\right).$
@@
$={1\over n}\sum_{i=1}^{n}\left[(X_{i}-{M}Y_{i})(X_{i}-{M}Y_{i})^{\top}-\Sigma_{w}\right].$
Recall that, conditioning ${\mathbf{Y}}$, $W_{i}=X_{i}-{M}Y_{i}$ for $i\in[n]$ are independent, $\mathcal{N}_{p}(0,\Sigma_{w})$. The result follows from standard concentration inequalities for the quadratic term of Gaussian random vectors. ∎
C.1 Concentration inequalities related with $\widehat{\Sigma}-\Sigma$
Recall that $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$ and the definition of the event $\mathcal{E}_{\pi}$ is given in (B.1).